In a cloud chamber experiment, a proton enters a uniform 0.340 T magnetic field directed perpendicular to its motion. You measure the proton's path on a photograph and find that it follows a circular arc of radius 6.30 cm .
Answer #1

B = 0.340 T
R = 0.063 m

When the proton follows a circular path, the magnetic force (F₁ = qvB) equates to centripetal force (F₂ = mv² / R) :
qvB = mv² / R
qB = mv / R
v = qBR / m
v = (1.6*10^-19 * 0.34 * 0.063)/ (1.67 *10^-27)
v = 2.05 * 10^6 m/s
Speed of proton, v = 2.05 * 10^6 m/s